The nuclei of the O2molecule are separated by 1.20 × 10–10m. The mass of each oxygen atom in the molecule is 2.66 × 10–26kg. (a) Determine the rotational energies of an oxygen molecule in electron volts for the levels corresponding to J = 0, 1, and 2.(b) The effective force constant k between the atoms in the oxygen molecule is 1177 N/m. Determine the vibrational energies (in electron volts) corresponding to v = 0, 1, and 2.


1
Expert's answer
2021-07-05T08:41:21-0400

(a) The reduced mass of the O2 is

"\mu= \frac{m_om_o}{m_o+m_o} \\

= \frac{16.00 \times 16.00}{16.00+16.00} = 8 \;u \\

= 8 \times 1.66 \times 10^{-27} = 1.33 \times 10^{-26} \; kg"

The momentum inertia is

"I = \mu r^2 \\

= 1.33 \times 10^{26} \times (1.20 \times 10^{-10})^2 \\

= 1.91 \times 10^{-46} \; kg \times m^2"

The rotational energies are

"E_{rot}= \frac{h^2}{2l}J(J+1) \\

= \frac{(6.626 \times 10^{-34 \;J \;s/2 \pi )^2}}{2 \times 1.91 \times 10^{-46} \;kg \; m^2}J(J+1) \\

= (2.91 \times 10^{-23}\;J) J(J+1)"

For J=0,1,2

"E_{rot}= 0, 3.63 \times 10^{-4}\;eV, 1.09 \times 10^{-3}\;eV"

(b) The vibrational energies are given by

"E_{vib}=(v+\frac{1}{2})h \sqrt{\frac{k}{\mu}} \\

=(v+ \frac{1}{2})(\frac{6.626 \times 10^{-34}J \;s}{2 \pi}) \sqrt{ \frac{1177 \;N/m}{8(1.66 \times 10^{-27} \;kg)} } \\

= (v+ \frac{1}{2})(3.14 \times 10^{-20}\;J)( \frac{1 \;eV}{1.602 \times 10^{-19} \;J} ) \\

= (v+ \frac{1}{2})(0.196 \;eV)"

For v=0,1,2

"E_{vib}=0.098 \;eV, 0.294 \;eV, 0.490 \;eV"